OpenStudy (anonymous):

how to solve : 2y = x + 10 : x + 2y - 6 = 0

OpenStudy (anonymous):

isolate x in the first equation to get x=2y-10 and substitute this into the second to get 2y-10+2y-6=0 which simplifies to 4y-16=0 4y=16 y=4 replace that into one of the equations, say the first? and you get 2(4)=x+10 8=x+10 8-10=x x=-2

OpenStudy (anonymous):

how to solve : 3x - y = -2 4x + 2y - 1 = 0

OpenStudy (anonymous):

Follow the exact same procedure, isolate a variable in one equation, plug it into the second and solve.

OpenStudy (anonymous):

how did you get -10 in the first equation ?

OpenStudy (anonymous):

By subtraction 10 from both sides...

OpenStudy (anonymous):

ok thanks